Exercise 9-6

Rottino Company purchased a new machine on October 1, 2015, at a cost of $121,800. The company estimated that the machine will have a salvage value of $10,400. The machine is expected to be used for 10,400 working hours during its 5-year life.

Compute the depreciation expense under straight-line method for 2015. (Round answer to 0 decimal places, e.g. 2,125.)

Compute the depreciation expense under units-of-activity for 2015, assuming machine usage was 1,500 hours. (Round per hour value to 2 decimal places, e.g. 21.25 and final answer to 0 decimal places, e.g. 2,125.)

Compute the depreciation expense under declining-balance using double the straight-line rate for 2015 and 2016. (Round answers to 0 decimal places, e.g. 2,125.)
Answer #1

a) 2015 Dep (Straight line) = (121800-10400/5)*3/12 = 5570

b) 2015 Dep (Unit of production) = (121800-10400)/10400 = 10.71*1500 = 16065

c) Double decline balance = 100/5*2 = 40%

2015 2016
Double decline balance 121800*40%*3/12 = 12180 121800*40%*9/12+121800*60%*40%*3/12 = 43848